# Dichotomy theorem separating complete integrability and non-integrability of isotropic spin chains

### v2

@@ Abstract
In contrast to integrability, theoretical aspects of non-integrability have not yet been fully addressed, though the absence of nontrivial local conserved quantities is a relevant property to various phenomena including thermalization Lan ; EF ; SM17 ; MS17 ; SMreply and response to external drives Kub ; SK ; SF ; Zot . Within the sparse literature devoted to non-integrability, one traditional approach is restricting the solution method to the Yang-Baxter equation and testing the validity of its lowest-order relation (the Reshetikin condition) KS82 ; Ken92 ; BY94 ; BN14 ; Hie92 ; LPR19 ; Lee20 ; Lee21 ; Mai24 ; LP24 . However, this approach suffers from an inherent limitation: we cannot rule out the possibility that a model is not solvable by the Yang-Baxter equation but is still solvable by another, possibly yet unknown, solution method. A more recent approach directly proves the absence of nontrivial local conserved quantities, which was invented by Ref Shi19 , and by applying this technique various quantum spin systems have been proven to be non-integrable Chi24-1 ; Shi24 ; PL24-1 ; CY24 ; ST24 ; Chi24-2 ; PL24-2 ; HYC24 ; YCS24-1 ; YCS24-2 ; Shi25 ; FT25 . Notably, symmetric $S=1/2$ systems in one YCS24-1 ; YCS24-2 ; Shi25 and higher dimensions ST24 ; Chi24-2 are completely classified into completely integrable and non-integrable. Another important progress is brought by Hokkyo Hok25 , which observes similarity among many previous non-integrability proofs and clarifies the sufficient conditions for proving non-integrability. Precisely, it is shown that there are two key operator equations and if these two operator equations respectively have specific forms of solutions as unique solutions, then the non-integrability proof surely works.
@@
Yang-Baxter equation and Reshetikin condition.— In order to explain the significance of our result, we here briefly introduce the Yang-Baxter equation and the Reshetikin condition (for details, see the Supplemental Material SM ). The Yang-Baxter equation is a relation on a one-parameter family of a matrix $R_{ij}(\lambda)$ on two sites $i$ and $j$ given by
@@
where $h$ is the local Hamiltonian and $\Pi$ is a swap operator. The third and higher-order terms are not universal. Plugging Eq. (6) into the Yang-Baxter equation (5) and compare the lowest nontrivial order, $\lambda\mu^{2}$, we obtain the Reshetikin condition claiming that $D$ in Eq. (2) can be expressed as KS82 ; Ken92
@@
Implications of the main theorem.— Remarkably, our theorem confirms that the Reshetikin condition does not miss integrable systems in isotropic spin chains, justifying the approach to non-integrability based on the Reshetikin condition explained in the introduction (i.e., asserting non-integrability by the violation of the Reshetikin condition) Ken92 ; BY94 . The integrable part ($D^{3}=0$) of our theorem contains the Reshetikin condition, since the Reshetikin condition (7) is equivalent to the condition $[Q^{3},H]=0$ with Eq. (4). Note that our dichotomy theorem imposes that if an infinite family of local conserved quantities exists, the Reshetikin condition should be satisfied. Hence, our theorem establishes that the Reshetikin condition is the necessary and sufficient condition for complete integrability in isotropic spin chains. In light of the fact that some integrable models including the Hubbard model do not satisfy the Reshetikin condition, it is striking that all isotropic integrable models admit the Reshetikin condition.
@@
This theorem also serves as the complete classification of complete integrability and non-integrability up to $S=13.5$. Kennedy Ken92 and Batchelor and Yung BY94 have examined all isotropic spin chains satisfying the Reshetikin condition less than or equal to $S=13.5$ by computer search and found that all the spin chains satisfying the Reshetikin condition are in fact exhausted by known integrable systems solvable by the Yang-Baxter equation. Precisely, there are four infinite sequences of integrable systems and one exceptional integrable system for $S=3$ (see End matter), which correspond to four infinite series of compact Lie groups and one exceptional Lie group satisfying SU(2) symmetry BY94 . Combining this fact and our theorem, we conclude that these four infinite sequences of integrable systems and one exceptional integrable system are indeed the complete list of integrable isotropic spin chains up to $S=13.5$.
@@
Proof for $D^{3}=0$.— We here consider the case of $D^{3}=0$. We here employ a recent result Hok25-2 stating that if the 3-local conserved quantity $Q^{3}$ given by Eq. (4) conserves ($[Q^{3},H]=0$), then all $k$-local conserved quantities $Q^{k}$ given by Eq. (3) also conserve ($[Q^{k},H]=0$) (see Lemma. S.1 in the Supplemental Material SM ). Hence, it suffices to show the presence of a 3-local conserved quantity, or the Reshetikin condition. By comparing the condition $D^{3}=0$ to the Reshetikin condition (7), our task boils down to that 2-support operators in $D$ take the form of $E_{i+1,i+2}-E_{i,i+1}$. We shall show this fact as follows.
@@
Our result shows that all integrable isotropic spin chains satisfy the lowest-order Yang-Baxter equation (the Reshetikin condition). Moreover, restricting isotropic spin chains with $S\leq 13.5$, all integrable isotropic spin chains satisfy the Yang-Baxter equation, meaning that the Yang-Baxter equation is the unique solution method in this regime. On the other hand, again the Hubbard model does not admit the original Yang-Baxter equation Lee21 . On the basis of the above, we expect that the Yang-Baxter equation is not always universal but indeed universal in some classes including isotropic spin chains. We note that the technique of non-integrability proofs is also extended to write down local conserved quantities of integrable systems NF20 ; YF23 . The non-integrability proof approach will pave the way to clarify how universally the Yang-Baxter equation applies, which merits further research.
@@ I.2 Reshetikin condition
Defining unknown 2-support operator $E:=h^{3}-6R^{(3)}$, we arrive at the Reshetikin condition s-KS82 ; s-Ken92
@@
The Reshetikin condition is thus the lowest-order nontrivial relation of the Yang-Baxter equation. This relation is sometimes employed to seek numerically new integrable models s-LMS25 .
@@ I.3 Boost operator
which is also equivalent to the Reshetikin condition (S.18).
If Eq. (S.28) (or the Reshetikin condition) holds, then $Q^{4}$ is also the 4-local conserved quantity, which is proven as

### v3

@@ Abstract
In contrast to integrability, theoretical aspects of non-integrability have not yet been fully addressed, though the absence of nontrivial local conserved quantities is a relevant property to various phenomena including thermalization Lan ; EF ; SM17 ; MS17 ; SMreply and response to external drives Kub ; SK ; SF ; Zot . Within the sparse literature devoted to non-integrability, one traditional approach is restricting the solution method to the Yang-Baxter equation and testing the validity of its lowest-order relation (the Reshetikhin condition) KS82 ; Ken92 ; BY94 ; BN14 ; Hie92 ; LPR19 ; Lee20 ; Lee21 ; Mai24 ; LP24 . However, this approach suffers from an inherent limitation: we cannot rule out the possibility that a model is not solvable by the Yang-Baxter equation but is still solvable by another, possibly yet unknown, solution method. A more recent approach directly proves the absence of nontrivial local conserved quantities, which was invented by Ref Shi19 , and by applying this technique various quantum spin systems have been proven to be non-integrable Chi24-1 ; Shi24 ; PL24-1 ; CY24 ; ST24 ; Chi24-2 ; PL24-2 ; HYC24 ; YCS24-1 ; YCS24-2 ; Shi25 ; FT25 . Notably, symmetric $S=1/2$ systems in one YCS24-1 ; YCS24-2 ; Shi25 and higher dimensions ST24 ; Chi24-2 are completely classified into completely integrable and non-integrable. Another important progress is brought by Hokkyo Hok25 , which observes similarity among many previous non-integrability proofs and clarifies the sufficient conditions for proving non-integrability. Precisely, it is shown that there are two key operator equations and if these two operator equations respectively have specific forms of solutions as unique solutions, then the non-integrability proof surely works.
@@
Yang-Baxter equation and Reshetikhin condition.— In order to explain the significance of our result, we here briefly introduce the Yang-Baxter equation and the Reshetikhin condition (for details, see the Supplemental Material SM ). The Yang-Baxter equation is a relation on a one-parameter family of a matrix $R_{ij}(\lambda)$ on two sites $i$ and $j$ given by
@@
where $h$ is the local Hamiltonian and $\Pi$ is a swap operator. The third and higher-order terms are not universal. Plugging Eq. (6) into the Yang-Baxter equation (5) and compare the lowest nontrivial order, $\lambda\mu^{2}$, we obtain the Reshetikhin condition claiming that $D$ in Eq. (2) can be expressed as KS82 ; Ken92
@@
Implications of the main theorem.— Remarkably, our theorem confirms that the Reshetikhin condition does not miss integrable systems in isotropic spin chains, justifying the approach to non-integrability based on the Reshetikhin condition explained in the introduction (i.e., asserting non-integrability by the violation of the Reshetikhin condition) Ken92 ; BY94 . The integrable part ($D^{3}=0$) of our theorem contains the Reshetikhin condition, since the Reshetikhin condition (7) is equivalent to the condition $[Q^{3},H]=0$ with Eq. (4). Note that our dichotomy theorem imposes that if an infinite family of local conserved quantities exists, the Reshetikhin condition should be satisfied. Hence, our theorem establishes that the Reshetikhin condition is the necessary and sufficient condition for complete integrability in isotropic spin chains. In light of the fact that some integrable models including the Hubbard model do not satisfy the Reshetikhin condition, it is striking that all isotropic integrable models admit the Reshetikhin condition.
@@
This theorem also serves as the complete classification of complete integrability and non-integrability up to $S=13.5$. Kennedy Ken92 and Batchelor and Yung BY94 have examined all isotropic spin chains satisfying the Reshetikhin condition less than or equal to $S=13.5$ by computer search and found that all the spin chains satisfying the Reshetikhin condition are in fact exhausted by known integrable systems solvable by the Yang-Baxter equation. Precisely, there are four infinite sequences of integrable systems and one exceptional integrable system for $S=3$ (see End matter), which correspond to four infinite series of compact Lie groups and one exceptional Lie group satisfying SU(2) symmetry BY94 . Combining this fact and our theorem, we conclude that these four infinite sequences of integrable systems and one exceptional integrable system are indeed the complete list of integrable isotropic spin chains up to $S=13.5$.
@@
Proof for $D^{3}=0$.— We here consider the case of $D^{3}=0$. We here employ a recent result Hok25-2 stating that if the 3-local conserved quantity $Q^{3}$ given by Eq. (4) conserves ($[Q^{3},H]=0$), then all $k$-local conserved quantities $Q^{k}$ given by Eq. (3) also conserve ($[Q^{k},H]=0$) (see Lemma. S.1 in the Supplemental Material SM ). Hence, it suffices to show the presence of a 3-local conserved quantity, or the Reshetikhin condition. By comparing the condition $D^{3}=0$ to the Reshetikhin condition (7), our task boils down to that 2-support operators in $D$ take the form of $E_{i+1,i+2}-E_{i,i+1}$. We shall show this fact as follows.
@@
Our result shows that all integrable isotropic spin chains satisfy the lowest-order Yang-Baxter equation (the Reshetikhin condition). Moreover, restricting isotropic spin chains with $S\leq 13.5$, all integrable isotropic spin chains satisfy the Yang-Baxter equation, meaning that the Yang-Baxter equation is the unique solution method in this regime. On the other hand, again the Hubbard model does not admit the original Yang-Baxter equation Lee21 . On the basis of the above, we expect that the Yang-Baxter equation is not always universal but indeed universal in some classes including isotropic spin chains. We note that the technique of non-integrability proofs is also extended to write down local conserved quantities of integrable systems NF20 ; YF23 . The non-integrability proof approach will pave the way to clarify how universally the Yang-Baxter equation applies, which merits further research.
@@ I.2 Reshetikhin condition
Defining unknown 2-support operator $E:=h^{3}-6R^{(3)}$, we arrive at the Reshetikhin condition s-KS82 ; s-Ken92
@@
The Reshetikhin condition is thus the lowest-order nontrivial relation of the Yang-Baxter equation. This relation is sometimes employed to seek numerically new integrable models s-LMS25 .
@@ I.3 Boost operator
which is also equivalent to the Reshetikhin condition (S.18).
If Eq. (S.28) (or the Reshetikhin condition) holds, then $Q^{4}$ is also the 4-local conserved quantity, which is proven as